OpenStudy (anonymous):

Could an elephant walking slowly across a field have more kinetic energy than a cheetah chasing its prey? Explain. I know that it is possible because the elephant has a greater mass than the cheetah, but the cheetah has a greater speed.. \(E_k=\frac{1}{2}mv^2\) so how to explain this properly to get a full mark?

OpenStudy (anonymous):

whatever you just said.. is right..

OpenStudy (anonymous):

KE = 1/2 mv^2 So velocity matters the most, but since the mass of the elephant is so much more, then the mass could beat the KE. Use the example that a cheetah weighs 45 kg and an elephant weighs at least 5000 kg So if the elephant was going 2 m/s the equation would be: KE = 1/2 (5000 kg) * (2 m/s)^2 = 10,000 Joules of energy If the cheetah weighs 45 kg but runs 65 mph ( about 29 m/s) then : KE = 1/2 (45 kg ) ( 29 m/s)^2 = 18,922 J Since the cheetah with it's high velocity, the elephant would have to either run faster or have a lot more mass such as: KE = 1/2 ( x kg ) (2 m/s) = 18,922 J solve x for needed kg weight of elephant : X = 18,922 kg......that would have to be one giant elephant to beat the KE of a cheetah So for the elephant to beat the cheetah it has to have a mass of about 19000 kg or else get faster

OpenStudy (anonymous):

Well, to put numbers to things.........an average cheetah weighs 72kg and can hit speeds of 33.3m/s (120 km/h), so such a cheetah would have a kinetic energy of 39.9kJ. An African Bush elephant can run up to about 40km/h (11.1m/s). Given those numbers, the elephant would only need to weigh 648kg. The real answer isn't straight forward because it depends on how fast and how big each of the animals is. My results and @biire2u's results defer, because of the assumptions each of us made. What that means--besides the fact that it "depends"--is that the elephant's mass needs to be as follows:\[m _{e}>m _{c}(\frac{ vc }{ ve })^{2}\]where vc is the cheetah's velocity and ve is the elephant's velocity.

OpenStudy (anonymous):

So mathematically speaking, the elephant's kinetic energy can exceed that of a cheetah.

OpenStudy (anonymous):

The question said " an elephant is walking slowly across the field". 2 m/s is about 6 mph, which is a fast walk to a human...and I assume a slow walk to an elephant

OpenStudy (anonymous):

You're right. I forgot that part. The question is, however, ill posed, so strictly speaking my result in terms of the equation I gave would be the best answer.

OpenStudy (anonymous):

So the normal answer is no, an elephant would not have enough kinetic energy to surpass a cheetah in full pursuit, if the elephant was walking slow. The elephant would have to be almost 42,000 lbs or more, and no elephant on record is close to that

OpenStudy (anonymous):

Actually, as I said, it's an ill posed question.

OpenStudy (anonymous):

Sorry for the redundancy..I was off working....where I'm supposed to be, but find this place more fun for some reason...

OpenStudy (anonymous):

hmm.. everyone has a point ^.^ this is fun! thanks for your awesome explanations :))